# All-in-one Graph-based Indexing for Hybrid Search on GPUs

### v1

@@ 3 $cand\leftarrow$ nodes containing user-specified entities;
7 $cand\leftarrow$ nodes with small vector length;
@@ 5.1. Experiment Settings
For comprehensive evaluations, we use 6 real-world datasets of varying scales, which have been widely used in related works (Zhang et al., 2024; Wang et al., 2025; Jimenez Gutierrez et al., 2024; Bruch et al., 2024a). Among them, NaturalQuestions (NQ(Wang et al., 2024b) and MS MARCO (MS(Nguyen et al., 2016) include simple queries, while 2WikiMultiHopQA (WM(Lv and Zhai, 2011) and HotpotQA (HP(Yang et al., 2018) contain complex, multi-hop queries. Table 1 summarizes the detailed information of each dataset. We employ the BGE-M3 model (Chen et al., 2024) to generate the dense vectors with a dimension of 1024, the SPLADE model (Formal et al., 2021) to generate the sparse vectors, and the BM25 algorithm (Robertson, 2025) to generate the full-text vectors.
@@
Figure 8. Comparison of all methods on 4 real-world datasets.
@@
We evaluate our proposed Allan-Poe against 6 state-of-the-art competitors representing both hybrid and single-path retrieval paradigms:
@@
  • Infinity (inf, 2025) is a modern database featuring with efficient hybrid search. It adopts the multi-path retrieval paradigm, which supports combinations of dense vector, sparse vector, and full-text search.
  • ThreeRouteGPU is our implemented GPU-based hybrid search method adopting the multi-path retrieval paradigm. It constructs separate CAGRA indexes for dense vectors, learned sparse vectors, and statistical sparse vectors (reducing dimension via JL transformation). Results retrieved from the three routes are then fused using fixed weights.
  • HippoRAG (Jimenez Gutierrez et al., 2024) is one of the state-of-the-art GraphRAG algorithms integrating knowledge graph and dense vectors to enhance the retrieval of relevant documents. Due to high knowledge graph construction costs, we compare it with All-Poe using logical edges only on the smaller datasets NQ-9633 and WM-6119.
  • Allan-Poe is our proposed method adopting the fusion retrieval paradigm. We denote different retrieval configurations as: dense for dense vectors only, sparse for sparse vectors only, full for full-text only, TwoPath for dense+sparse combination, and ThreePath for all three paths combined.
@@
We evaluate the indexing efficiency by measuring the construction time, the retrieval efficiency by Queries Per Second (QPS), and the retrieval accuracy by nDCG@$k$. Without additional explanation, the value of $k$ is set to 10.
@@ 5.2. Overall Performance for Query Processing
In this section, we compare the QPS and nDCG@10 across all methods on four real-world datasets. All the compared hybrid search methods employ equal weighting for all retrieval paths unless specified otherwise. The CPU-based methods, i.e., SEISMIC and Infinity, utilize 48 threads during query processing. The results are depicted in Figure 8. For Infinity and SEISMIC, we report single data points since nDCG values show minimal variation with parameter tuning.
[…]
However, as noted in Section 2.2, additional retrieval paths do not guarantee improved accuracy in all scenarios. For instance, on datasets NQ and MS, Allan-Poe-TwoPath and Allan-Poe-sparse achieve optimal performance, respectively. This occurs because all participating paths influence final accuracy—on NQ and MS, while dense and sparse retrieval perform well, full-text retrieval underperforms and reduces overall accuracy. Moreover, for dataset MS, Allan-Poe-sparse using a single retrieval path outperforms the hybrid retrieval methods because all the documents in MS have a short length with simple semantics, which can be efficiently handled by sparse vectors. These results emphasize the importance of supporting flexible path combinations without index reconstruction, corresponding to the flexibility dimension in Figure 5.
5.3. Evaluations of Indexing Efficiency
@@
Allan-Poe 40.08 36.02 19.50 24.25 186 186 78 95
5.4. Evaluations of Keyword Retrieval
@@
To evaluate the effectiveness of keyword specification in queries, we employ the Qwen3 LLM (Yang et al., 2025a) to simulate users specifying required keywords for the first 100 queries from the representative datasets NQ and WM. Figure 9 compares results with and without keyword supplementation, reporting the highest nDCG@10 and corresponding QPS for each method. Keyword supplementation improves nDCG@10 by 1.2% on average across retrieval paths, with only a 3.2% QPS reduction, demonstrating its effectiveness. However, some methods show minimal accuracy gains (e.g., 0.2% for three-path retrieval on NQ), as many relevant documents already contain the required keywords, limiting the filtering impact. Interestingly, certain retrieval paths with keyword supplementation achieve higher QPS than their non-supplemented counterparts (e.g., full-text search on NQ and dense vector search on WM). This suggests that keyword constraints enable earlier convergence to optimal accuracy by focusing the search on more relevant nodes within a small candidate pool in query process.
@@ 5.5. Evaluations of Logical Augmentation
Table 3. Comparison of retrieval methods on small datasets. ‘kg’ denotes the methods augmented by the knowledge graph.
@@
Allan-Poe-ThreePath (kg) 0.751 8520.28 0.834 4126.26
To evaluate knowledge graph augmentation while managing the knowledge graph construction costs, we use Qwen3 to construct knowledge graphs for the two smaller datasets NQ-9633 and WM-6119, and simulate user queries by specifying required entities for the first 100 queries. Table 3 presents the evaluation results. Overall, Allan-Poe-ThreePath (kg) exhibits the highest nDCG@10 while maintaining competitive QPS. On NQ-9633, which contains simple queries without multi-hop reasoning, knowledge graph augmentation provides modest accuracy improvements—particularly for single-path Allan-Poe variants, with slight gains for the three-path configuration. In contrast, the accuracy improvements on WM-6119 are significant due to the complex multi-hop queries in this dataset, demonstrating the effectiveness of the knowledge graph integration in Allan-Poe. While HippoRAG (a GraphRAG method) achieves higher nDCG@10 than approaches without knowledge graphs, it still underperforms compared to Allan-Poe. This gap occurs because HippoRAG does not effectively integrate knowledge graph information with document-level semantic similarity, and its community search can introduce redundant documents that impair query efficiency despite being useful for global queries. Consequently, integrating knowledge graphs with document-level vector search represents a promising direction for future research.
@@ 5.7. Weights of Retrieval Paths
To investigate the impact of retrieval path weighting, we evaluate Allan-Poe-TwoPath and Allan-Poe-ThreePath under various weight configurations. For the two-path configuration (dense + sparse vectors), the fused distance is computed as $\alpha\cdot\text{sim}_{d}(q,d)+(1-\alpha)\cdot\text{sim}_{s}(q,d)$, where $\alpha\in[0,1]$, and $q$, $d$ denote query and document respectively. For the three-path configuration, the distance function is $\alpha\cdot[\text{sim}_{d}(q,d)+w_{\text{opt}}\cdot\text{sim}_{s}(q,d)]+(1-\alpha)\cdot\text{sim}_{f}(q,d)$, where $w_{\text{opt}}$ represents the optimal dense-sparse weight derived from Allan-Poe-TwoPath evaluations. Results are presented in Figure 12.
@@ 5.8. Evaluations of Data Insertion
As established in Section 4.1, Allan-Poe supports efficient data insertion and mark-and-delete operations to accommodate data updates. We evaluate both insertion efficiency and its impact on retrieval quality by inserting varying data volumes into pre-built hybrid indexes and measuring subsequent search performance. As shown in Figure 13, the performance decrease of the updated index is marginal compared to the rebuilt index, which is up to 1% of nDCG@10 with the same QPS. Figure 13 shows that the updated index experiences only marginal performance degradation compared to a complete rebuild, with at most a 1% reduction in nDCG@10 while maintaining equivalent QPS. Furthermore, as shown in Table 2, our insertion strategy incorporates 20% new data with only 14.5% of the computational overhead required for a full index rebuild. These results demonstrate Allan-Poe’s capability to efficiently handle data updates in dynamic environments.
@@
Table 4. Comparison of indexing overhead.
Datasets Rebuild Insert 20% Insert 10% Insert 5%
NQ 40.08s 5.82s 2.86s 1.40s
WM 19.50s 2.20s 1.06s 0.51s
@@ 6. Conclusion
This paper presents Allan-Poe, a unified, GPU-accelerated hybrid index that integrates dense vector, sparse vector, full-text, and knowledge graph retrieval. We first analyze the limitations of existing retrieval paradigms and derive design principles for effective hybrid indexing. Guided by these principles, we design an all-in-one graph-based index featuring an isolated heterogeneous edge storage that integrates multiple retrieval paths within a unified structure while minimizing maintenance overhead. Furthermore, we optimize index construction through hybrid distance acceleration, RNG-IP joint pruning, and keyword-aware neighbor recycling, leveraging massive GPU parallelism to accelerate the entire construction pipeline. Finally, we introduce a unified query processing strategy that dynamically fuses information from all retrieval paths to achieve high-accuracy results. Our approach also innovatively augments document-level vector search with knowledge graph reasoning to handle complex queries. Comprehensive experiments on real-world datasets demonstrate that Allan-Poe consistently outperforms state-of-the-art methods in both efficiency and accuracy.

### v2

@@ 3 $cand\leftarrow$ nodes containing user-specified entities;
7 $cand\leftarrow$ nodes with large vector length;
@@ 28       $cand\leftarrow cand\cup\{o\}$, push $o$ to topktopk;
4.2. Flexible Query Processing on GPU
4.2.1. Query on the Base Semantic Edges.
As established in Section 3.2, base semantic edges constructed from fused vectors support arbitrary weight combinations across retrieval paths. Given a weight vector $w=[w_{d},w_{s},w_{f}]$ for dense, learned sparse, and lexical sparse vectors respectively, the query vector is formulated as $f_{\text{concat}}(q,w_{d},w_{s},w_{f})=[w_{d}*f_{\text{dense}}(q),w_{s}*f_{\text{sparse}}(q),w_{f}*f_{\text{full}}(q)]$. Single-path retrieval is achieved by setting the corresponding weight to 1 (or any non-zero value) and others to 0. For instance, the fused query vector $f_{concat}(q,1,0,0)$ retrieves documents using only dense vector similarity. This approach extends naturally to two-path and three-path configurations. To optimize search efficiency, we select entry points from nodes with the largest vector norms. The computationally intensive hybrid distance calculations during query processing are accelerated using the same GPU-optimized strategy described in Section 4.1. The only difference is that during the set intersection operation, each thread fetches an index from the document’s sparse vector and searches for it in the query’s sparse vector. This design reduces time complexity by searching the typically smaller query vector.
4.2.2. Query with Keyword Augmentation.
@@ 5.1. Experiment Settings
For comprehensive evaluations, we use 6 real-world datasets of varying scales, which have been widely used in related works (Zhang et al., 2024; Wang et al., 2025; Jimenez Gutierrez et al., 2024; Bruch et al., 2024a). Among them, NaturalQuestions (NQ(Wang et al., 2024b) and MS MARCO (MS(Nguyen et al., 2016) include simple queries, while 2WikiMultiHopQA (WM(Lv and Zhai, 2011) and HotpotQA (HP(Yang et al., 2018) contain complex, multi-hop queries. Table 1 summarizes the detailed information of each dataset. We employ the BGE-M3 model (Chen et al., 2024) to generate the dense vectors with a dimension of 1024, the SPLADE model (Formal et al., 2021) to generate the sparse vectors, and the BM25 algorithm (Robertson, 2025) to generate the full-text vectors. Considering the high dimensionality of various types of vectors, we extract at most one million documents from each dataset, which aligns with the typical data segment size in vector databases (Wang et al., 2024c, 2021b).
@@
Figure 8. Comparison of all methods on 4 real-world datasets. Paths in hybrid search methods share equal weights.
@@
We evaluate Allan-Poe against 6 state-of-the-art competitors with both hybrid and single-path retrieval paradigms:
@@
  • Infinity (inf, 2025; Wang et al., 2025) is a modern database featuring with efficient hybrid search. It adopts the multi-path retrieval paradigm, which supports combinations of dense vector, sparse vector, and full-text search.
  • ThreeRouteGPU is our implemented GPU-based hybrid search method adopting the multi-path retrieval paradigm. It constructs separate CAGRA indexes for dense vectors, learned sparse vectors, and lexical sparse vectors (reducing dimension via JL transformation). Results retrieved from the three routes are then fused using fixed weights.
  • HippoRAG (Gutiérrez et al., 2025) is a state-of-the-art GraphRAG algorithm. Following evaluations in recent studies (Zhou et al., 2026; Jimenez Gutierrez et al., 2024), we evaluate knowledge graph augmentation on two benchmark datasets, NQ-9633 (50k entities, 68k relations) and WM-6119 (30k entities, 31k relations).
  • Allan-Poe is our proposed hybrid search method, with configurations denoted as: dense (dense vectors only), sparse (sparse vectors only), full (full-text only), TwoPath (dense and sparse vectors), and ThreePath (all three retrieval paths combined). All configurations leverage the same hybrid index without reconstruction.
@@
We evaluate the indexing efficiency by measuring the construction time, the retrieval efficiency by Queries Per Second (QPS), and the retrieval accuracy by nDCG@$k$ (higher is better). Without additional explanation, the value of $k$ is set to 10.
@@ 5.2. Overall Performance for Query Processing
In this section, we compare the QPS and nDCG@10 across all methods on real-world datasets. All hybrid search methods employ equal weighting for all retrieval paths unless specified otherwise. The CPU-based methods, i.e., SEISMIC and Infinity, utilize 48 threads during query processing. The results are depicted in Figure 8. For Infinity and SEISMIC, we report single data points since nDCG values show minimal variation with parameter tuning (e.g., $k^{\prime}$).
@@
Allan-Poe 45.83 40.63 21.86 26.36 186 186 78 95
However, as noted in Section 2.2, additional retrieval paths do not guarantee improved accuracy in all scenarios. For instance, on datasets NQ and MS, Allan-Poe-TwoPath and Allan-Poe-sparse achieve optimal performance, respectively. This occurs because all participating paths influence the final accuracy—on NQ and MS. While dense and sparse retrieval perform well, full-text retrieval underperforms and reduces overall accuracy. Moreover, for dataset MS, Allan-Poe-sparse using a single retrieval path outperforms the hybrid retrieval methods because all the documents in MS have a short length with simple semantics, which can be efficiently handled by sparse vectors. These results emphasize the importance of supporting flexible path combinations without index reconstruction, corresponding to the flexibility dimension in Figure 5.
@@
5.3. Evaluations of Indexing Efficiency
[…]
5.4. Evaluations of Keyword Retrieval
To evaluate the effectiveness of keyword specification in queries, we employ the Qwen3 LLM (Yang et al., 2025a) to simulate users specifying required keywords for the first 100 queries from the representative datasets NQ and WM. Figure 9 compares results with and without keyword supplementation, reporting the highest nDCG@10 and corresponding QPS for each method. Keyword supplementation improves nDCG@10 by 1.2% on average across retrieval paths, with only a 3.2% QPS reduction, demonstrating its effectiveness. However, some methods show minimal accuracy gains (e.g., 0.2% for three-path retrieval on NQ), as many relevant documents already contain the required keywords, limiting the filtering impact. Interestingly, certain retrieval paths with keyword supplementation achieve higher QPS than their non-supplemented counterparts (e.g., full-text search on NQ and dense vector search on WM). This suggests that keyword constraints enable earlier convergence to optimal accuracy by focusing the search on more relevant nodes within a small candidate pool in the query process.
@@ 5.5. Evaluations of Logical Augmentation
Table 3. Comparison of retrieval methods on small datasets. ‘KG’ denotes the methods augmented by knowledge graphs.
@@
Allan-Poe-ThreePath (+KG) 0.751 8520.28 0.834 4126.26
To evaluate knowledge graph augmentation while managing the knowledge graph construction costs, we use Qwen3 to construct knowledge graphs for the two benchmark datasets NQ-9633 and WM-6119, and simulate user queries by specifying required entities for the first 100 queries. As mentioned in Section 3.3, logical edge augmentation is a trade-off between improvements in accuracy and costs of constructing the knowledge graph. Applying it to small-scale, high-value data is a cost-effective option. Table 3 presents the evaluation results. Overall, Allan-Poe-ThreePath (+KG) exhibits the highest nDCG@10 while maintaining competitive QPS. On NQ-9633, which contains simple queries without multi-hop reasoning, knowledge graph augmentation provides modest accuracy improvements—particularly for single-path Allan-Poe variants, with slight gains for the three-path configuration. Conversely, the accuracy improvements on WM-6119 are significant due to the complex multi-hop queries in this dataset. While HippoRAG (a GraphRAG method) achieves higher nDCG@10 than approaches without knowledge graphs, it still underperforms compared to Allan-Poe. This gap occurs because HippoRAG does not effectively integrate knowledge graph information with document-level semantic similarity, and its community search can introduce redundant documents that impair query efficiency despite being useful for global queries.
To evaluate the effectiveness of logical edge augmentation at scale, we assess Allan-Poe with and without knowledge graph enhancement on the larger NQ and WM datasets. The results, summarized in Table 4, show consistent accuracy gains with knowledge graph integration. Notably, improvements are more pronounced on the WM dataset, which contains complex, multi-hop queries. These findings underscore the significant potential of enhancing document-level semantic search with fine-grained, entity-level logical structures from knowledge graphs.
Table 4. Effectiveness of KG augmentation on large datasets.
Methods NQ WM
nDCG@10 QPS nDCG@10 QPS
Allan-Poe-ThreePath 0.580 28952.61 0.649 28335.84
Allan-Poe-ThreePath (+KG) 0.631 26503.97 0.769 25010.07
@@ 5.6. Effectiveness of Heterogeneous Edges
In this subsection, we evaluate the effectiveness of the three types of heterogeneous edges in our hybrid graph-based index, which are established in Section 4.1.
@@ 5.7. Weights of Retrieval Paths
To investigate the impact of retrieval path weighting and the robustness of the hybrid index for arbitrary weights, we evaluate Allan-Poe-TwoPath and Allan-Poe-ThreePath under various weight configurations. For the two-path configuration (dense + sparse vectors), the fused distance is computed as $\alpha\cdot\text{sim}_{d}(q,d)+(1-\alpha)\cdot\text{sim}_{s}(q,d)$, where $\alpha\in[0,1]$, and $q$, $d$ denote query and document respectively. For the three-path configuration, the distance function is $\alpha\cdot[\text{sim}_{d}(q,d)+w_{\text{opt}}\cdot\text{sim}_{s}(q,d)]+(1-\alpha)\cdot\text{sim}_{f}(q,d)$, where $w_{\text{opt}}$ represents the optimal dense-sparse weight derived from Allan-Poe-TwoPath evaluations. Results are presented in Figure 12.
@@ 5.8. Evaluations of Data Insertion
As established in Section 4.1, Allan-Poe supports efficient data insertion to accommodate data updates. We evaluate both insertion efficiency and its impact on retrieval quality by inserting varying data volumes into pre-built hybrid indexes and measuring subsequent search performance. As shown in Table 5, our insertion strategy incorporates 20% new data with only 13.7% of the computational overhead required for a full index rebuild. Furthermore, Figure 13 shows that the updated index experiences only marginal performance degradation compared to a complete rebuild, with at most 1% reduction in nDCG@10 while maintaining equivalent QPS. These results demonstrate Allan-Poe’s capability to handle data updates in dynamic environments efficiently.
Table 5. Comparison of indexing overhead.
Datasets Rebuild Insert 20% Insert 10% Insert 5%
NQ 45.83s 6.97s 3.43s 1.69s
WM 21.86s 2.67s 1.30s 0.64s
@@
5.9. Scalability
To assess the scalability of Allan-Poe, we conduct experiments on the MS dataset (~8.8M documents in total), scaling from 2 to 8 million documents using an NVIDIA A100 (40GB) GPU. As shown in Figure 14, both index size and construction time scale nearly linearly with dataset size. Notably, query latency exhibits a more favorable sub-linear (logarithmic) growth. These results demonstrate Allan-Poe’s robust scalability for both indexing and search. For datasets where the hybrid vectors exceed single-GPU memory, the graph-based architecture of Allan-Poe remains compatible with established horizontal and vertical scaling strategies (Li et al., 2025b; Wang et al., 2024c, 2021b), ensuring practical applicability to very large corpora.
(a) Index size.
(b) Index build time.
(c) Latency at nDCG of 0.55.
Figure 14. Impact of varying data size on dataset MS.
@@ 6. Conclusion
This paper presents Allan-Poe, a unified, GPU-accelerated hybrid index integrating dense vector, sparse vector, and full-text retrieval. We derive design principles for hybrid indexing and build an all-in-one graph-based index with isolated heterogeneous edge storage, enabling flexible path combinations with low overhead. The construction pipeline is optimized via hybrid distance computation, RNG-IP joint pruning, and keyword-aware recycling, fully leveraging GPU parallelism. At query time, we propose a dynamic fusion framework that combines all retrieval paths and incorporates knowledge graph structures to handle complex queries. Extensive experiments show that Allan-Poe outperforms state-of-the-art methods in both accuracy and efficiency.